# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 12.6 Banu Bedestenci Sönmez Liver Paired Exchange System
In June 2022, we received the initial dataset from Malatya and began running the system.135135135Before the deployment of the computerized system in June 2022, the Malatya team conducted 2-way exchanges on an emergency basis on two occasions: first in October 2021 and second in January 2022. Just days before my first visit to Turkey in three years, during the first week of July 2022, we surprised the Malatya team with an unexpected outcome that no one had anticipated, especially not at such an early stage.
@@ 12.6.1 Breaking New Ground: The First-Ever 4-Way Liver Exchange
After several entries into the liver exchange pool, the number of patient-donor pairs reached 12.136136136Although I use the phrase “patient-donor pairs,” it is not uncommon for patients to have multiple willing donors. Therefore, the intended meaning is patients with one or more donors. Upon receiving a new set of data, I couldn’t believe my eyes. As always, I reviewed it before Ünver ran the software he’d coded for the system. Immediately, I spotted a 4-way exchange alongside a separate 2-way exchange. Remarkably, it was already possible to match at least half of the patients in the pool! By then, even a 3-way liver exchange was unprecedented.137137137As it turns out, with the support of Alex Chan—a graduate student at Stanford Business School, a world-first 3-way liver exchange was carried out in Pakistan a few months earlier on March 17, 2022. We became aware of this development in December 2022 when its report became available online in JAMA Surgery (Salman et al., 2023).
@@
Once Ünver ran the software, he confirmed that the matching I identified is one of several “maximum-size” matchings. Although the largest exchange was a 5-way, an exchange the Malatya team was capable of implementing, performing a 4-way exchange along with a 2-way exchange would save one additional life. Conversely, to maximize the number of transplants, it was necessary to carry out a 4-way exchange and a 2-way exchange. With only 2-way and 3-way exchanges, one fewer patient, and with only 2-way exchanges, two fewer patients would receive transplants (Yilmaz et al., 2023). We immediately surprised Yılmaz with the good news.
@@
Even though it was quite late in Turkey, around 9 PM, when we shared the news, Yılmaz and much of his team were still at the hospital. Though we didn’t know it then, over time, we came to realize that this level of commitment is the norm in the Malatya team. In a public hospital, situated in one of the most isolated and often neglected parts of the country, Yılmaz not only built a highly skilled team but also an exceptionally dedicated one.
@@
Yılmaz promptly assembled his team, and by midnight, they had selected the exchanges they planned to carry out within a few days: a 4-way and a 2-way. The next morning, after informing the patients and their donors and confirming their continued interest in participating in the system, the Malatya team scheduled the 4-way exchange for July 5, 2022, and the 2-way exchange for July 7, 2022.
@@
The Malatya team successfully conducted a world-first 4-way liver exchange on its scheduled day, just one day before I flew to Turkey. Upon reaching Ankara the next day, I dropped off my son, Alp Derin, at my parents’ house and immediately flew to Malatya. On July 7th, 2022, I personally witnessed the operations for the week’s second liver exchange—this time a 2-way exchange. I also visited the patients and donors recovering from their surgeries for the 4-way exchange conducted two days earlier.
@@ 12.6.2 Heightened Importance of Size Incompatibility in Liver Exchange
The BBS-LPE System effectively transformed individual misfortunes into collective opportunities. The size incompatibility between the Afghan refugee and his pediatric patient, typically a source of despair, became a beacon of hope for others. In addition to this pediatric patient, three adult patients also received liver transplants: one faced blood type incompatibility with their registered donor, while the other two had donors with insufficiently sized right lobes. Notably, among the four pairs, only one encountered blood type incompatibility, while the remaining three faced size incompatibility.
When we asked Yılmaz how often prospective donors are unable to donate their left lobes, we received another surprising answer. He told us that it is extremely common. Under their transplantation norms, no more than 30% of donors who could otherwise safely donate a lobe could donate their left lobes! This information revealed another fundamental aspect of liver exchange. Donors who could safely donate their left lobes had very high value for the system because they are the only ones who can donate to pediatric patients.
@@ 12.6.3 The Pilot Phase in the First Year
For the year leading up to August 2023, we were disappointed that the system generated only a few additional exchanges, with numbers much lower than we had hoped. On the bright side, these exchanges included a 3-way exchange in April 2023, marking the third such exchange globally. This achievement followed the world’s first 3-way exchange conducted in Pakistan in March 2022 (Salman et al., 2023), which occurred a few months after our earlier 4-way exchange, and a second 3-way exchange in India in December 2022 (Soin et al., 2023). Including the two 2-way exchanges performed before the computerized system was deployed in June 2022, the Malatya team had conducted a total of 15 liver exchange transplants by August 2023.
While the contribution of liver exchange to the number of LDLTs performed by the Malatya team was much higher than that of any other liver exchange system globally up until that time, it fell short of our expectations. For example, the 10 transplants generated by the BBS-LPE System in 2022 corresponded to 4.3% of all LDLTs performed by the Malatya team. We had been hoping for at least 20%. Several reasons account for this relatively slow start.
@@
Yılmaz must be one of the most perceptive and hardest-working individuals I have ever known. Despite being devastated by the situation, he navigated through this challenging period by collaborating with us to document our experience with the BBS-LPE System for scholarly publication, with a specific focus on our groundbreaking 4-way exchange.
@@
Fortunately, the editorial process in medicine moves much faster than in economics. The timely online publication of Yilmaz et al. (2023) in July 2023 not only provided an opportunity to introduce the BBS-LPE System to the general Turkish public but also reversed the unfortunate dynamics that had gripped the entire Malatya team.
12.6.4 Milestones Post-Pilot Phase: World’s First 5-Way, 6-Way, and 7-Way Exchanges
In an effort to propel the Liver Transplantation Institute forward, İnönü University leadership seized the opportunity presented by the publication of Yilmaz et al. (2023) and hosted a media event on July 28, 2023. Participants from both the 4-way and 3-way exchanges attended the event, expressing praise for the BBS-LPE System and sharing their experiences with the wider public.
This unexpected accomplishment, emerging from a region still grappling with the aftermath of a once-in-a-century catastrophe, quickly made headlines across Turkey. Within weeks, a surge in patient participation became evident, leading to a remarkable increase in the number of transplants facilitated by the system—an outcome none of us could have anticipated.
@@ 12.6.4 Milestones Post-Pilot Phase: World’s First 5-Way, 6-Way, and 7-Way Exchanges
The public announcement marked the end of the BBS-LPE System’s pilot phase. Over the next seventeen months, it facilitated an additional 183 LDLTs—with an average of more than ten transplants per month—and brought the total to 198 by the end of 2024. This pace surpassed the outcomes of even the most active liver exchange programs by an order of magnitude. Before the BBS-LPE System’s introduction, only one center—Medanta in India—had performed as many as ten liver exchange transplants in a single calendar year. In contrast, this period witnessed a series of unprecedented milestones:
@@
  • World’s First Multi-Way Exchanges: The Malatya team performed the world’s first 5-way, 6-way, and 7-way exchanges in October 2023, January 2024, and July 2024, respectively (see Figure 9 for a depiction of the first 7-way liver exchange).
@@
  • Record-Breaking Overall Volume: By March 2024—less than two years after the system’s June 2022 launch—the Malatya team became the first in the world to surpass 100 cumulative liver exchange transplants.
@@
  • Unmatched Annual Performance: In October 2024, the team became the first to exceed 100 liver exchange transplants within a single calendar year.
As of December 2024, the BBS-LPE System accounts for both 7-way exchanges, all four 6-way, all four 5-way, and all eight 4-way conducted globally, along with eighteen of the twenty-one 3-way liver exchanges ever reported, firmly establishing itself as the global leader in liver exchange.
@@ 12.6.5 Comparison with Other Prominent Liver Exchange Programs
The BBS-LPE program, though not the first of its kind, has rapidly emerged as the global leader in liver exchange in just two years since its deployment. To understand the key factors behind this success, it is helpful to compare its outcomes with those of other prominent liver exchange programs worldwide.
In 2023, despite a catastrophic earthquake that disrupted life across the region for months, the BBS-LPE System facilitated 64 LDLTs, accounting for 27.7% of the 231 LDLTs performed by the Malatya team that year. Building on this achievement, the system enabled 122 LDLTs in 2024—an unprecedented 45.2% of the 270 LDLTs performed during this period.
Cumulatively, the system facilitated 2 liver exchange transplants in 2021, 10 in 2022, 64 in 2023, and 122 in 2024, bringing the total to 198 by the end of 2024. To put this into perspective, the global total of reported liver exchange transplants between 2002 and 2022 was fewer than 250 prior to the BBS-LPE System’s deployment. Remarkably, in just two years, the Malatya team facilitated approximately 80% of the global total of reported liver exchange transplants accumulated over the two decades preceding its launch.
Prior to the BBS-LPE system, the global leader in liver exchange was Medanta in India. Over thirteen years (2009 to 2022), this center conducted 91 liver exchange transplants, including one 3-way exchange and forty-four 2-way exchanges (Soin et al., 2023).
Two other major liver exchange programs are located at ASAN Medical Center in South Korea and Max Saket Hospital in India. Although the ASAN team performed the world’s first liver exchange in 2002, the number of liver exchange transplants in South Korea remained at 52 from 2002 to 2018, representing merely 0.4% of the LDLTs conducted in the country during this period (Kim, 2022). On the other hand, Agrawal et al. (2022) reported 34 liver exchange transplants at Max Saket Hospital from 2012 to 2021, constituting 1.45% of the LDLTs performed in the program during that timeframe.
@@
Notably, all liver exchanges at Max Saket Hospital and a significant majority at ASAN Medical Center involve 2-way exchanges between two blood-type incompatible pairs: one pair with a blood-type A patient and a blood-type B donor, and another pair with a blood-type B patient and a blood-type A donor. This limitation is likely the most significant factor explaining the relatively smaller number of liver exchange transplants performed at these centers.
@@
In the U.S., liver exchange programs are currently in their early phases. As of December 2024, a total of 173 liver exchange transplants have been reported in the U.S., an all-time high of 50 of them occurring in 2023 (OPTN, 2024a). Most activity is concentrated at two centers. The liver exchange program at the University of Texas Health San Antonio reported 23 liver exchange transplants from 2019 through October 2023 (Klair et al., 2024). These transplants were conducted through four 2-way exchanges, six 2-way non-directed donor chains, and one 3-way non-directed donor chain. Similarly, the program at the University of Pittsburgh Medical Center reported 20 liver exchange transplants from January 2019 through July 2021 (Gunabushanam et al., 2022), achieved through three 2-way exchanges and seven 2-way non-directed donor chains.

### v4

@@ 12.6 Banu Bedestenci Sönmez Liver Paired Exchange System
In June 2022, we received the initial dataset from Malatya and began running the system.150150150Before the deployment of the computerized system in June 2022, the Malatya team conducted 2-way exchanges on an emergency basis on two occasions: first in October 2021 and second in January 2022. Just days before my first visit to Turkey in three years, during the first week of July 2022, we surprised the Malatya team with an unexpected outcome that no one had anticipated, especially not at such an early stage.
@@ 12.6.1 Breaking New Ground: The First-Ever 4-Way Liver Exchange
After several entries into the liver exchange pool, the number of patient–donor pairs reached 12.151151151Although I use the phrase “patient–donor pairs,” it is not uncommon for patients to have multiple willing donors. Therefore, the intended meaning is patients with one or more donors. Upon receiving a new set of data, I couldn’t believe my eyes. As always, I reviewed it before Ünver ran the software he’d coded for the system. Immediately, I spotted a 4-way exchange alongside a separate 2-way exchange. Remarkably, it was already possible to match half of the patients in the pool! By then, even a 3-way liver exchange was unprecedented.152152152As it turns out, with the support of Alex Chan—now an Assistant Professor at Harvard Business School, and a graduate student at Stanford Business School at the time—a world-first 3-way liver exchange was carried out in Pakistan a few months earlier on March 17, 2022. We became aware of this development in December 2022, when its report appeared online in JAMA Surgery (Salman et al., 2023).
@@
Once Ünver ran the software, he confirmed that the matching I had identified was one of several “maximum-size” matchings. Although the largest possible exchange was a 5-way—an exchange the Malatya team was capable of implementing—performing a 4-way exchange together with a 2-way exchange would actually save one additional life. In fact, maximizing the number of transplants required exactly this combination of a 4-way and a 2-way exchange. Relying only on 2-way and 3-way exchanges would result in one fewer patient being transplanted, and restricting to 2-way exchanges alone would mean two fewer patients. (See Figure 32 for feasible transplants and liver exchanges during the first week of July 2022.) We immediately surprised Yılmaz with the good news.
Figure 32: Possible transplants and liver exchanges under the Banu Bedestenci Sönmez Liver Paired Exchange System at the Liver Transplantation Institute, İnönü University, Malatya, Turkey, during the first week of July 2022. Each node in the graphs indicates a patient–donor pair. Abbreviation: LPE: Liver paired exchange. Reproduced from Yilmaz et al. (2023), American Journal of Transplantation, licensed under $C$ BY-NC-ND 4.0.
@@
Even though it was quite late in Turkey—around 9 PM—when we shared the news, Yılmaz and much of his team were still at the hospital. Although we didn’t realize it at the time, it later became clear that this level of commitment is the norm for the Malatya team. In a public hospital located in one of the most isolated and often neglected parts of the country, Yılmaz built not only a highly skilled team but also an exceptionally dedicated one.
@@
Yılmaz promptly assembled his team, and by midnight, they had selected the exchanges they planned to carry out within a few days: the 4-way and 2-way exchanges that made the maximal size matching. The next morning, after informing the patients and their donors and confirming their continued interest in participating in the system, the Malatya team scheduled the 4-way exchange for July 5, 2022, and the 2-way exchange for July 7, 2022.
@@
The Malatya team successfully conducted a world-first 4-way liver exchange on its scheduled day, just one day before I flew to Turkey. As soon as I reached Ankara the next day, I dropped off my son, Alp Derin, at my parents’ house and immediately flew to Malatya. On July 7th, 2022, I personally witnessed the operations for the week’s second liver exchange—this time a 2-way exchange. I also visited the patients and donors recovering from their surgeries for the 4-way exchange conducted two days earlier.
@@ 12.6.2 Heightened Importance of Size Incompatibility in Liver Exchange
The BBS–LPE System effectively transformed individual misfortunes into collective opportunities. The size incompatibility between the Afghan refugee and his pediatric patient, typically a source of despair, became a beacon of hope for others. In addition to this pediatric patient, three adult patients also received liver transplants: one faced blood-type incompatibility with their registered donor, while the other two had donors with insufficiently sized right lobes. Notably, among the four pairs, only one encountered blood-type incompatibility, while the remaining three faced size incompatibility.
When we asked Yılmaz how often prospective donors are unable to donate their left lobes, we received another surprising answer: it is extremely common. According to their transplantation norms, no more than 30% of donors who could safely donate a lobe were eligible to donate their left lobes! This information revealed another fundamental aspect of liver exchange. Donors who could safely donate their left lobes had very high value for the system because they are the only ones who can donate to pediatric patients.
@@ 12.6.3 The Pilot Phase in the First Year
For the year leading up to August 2023, we were disappointed that the system generated only a few additional exchanges, with numbers much lower than we had hoped. On the bright side, these included a 3-way exchange in April 2023, marking the third such exchange globally. This achievement followed the world’s first 3-way exchange conducted in Pakistan in March 2022 (Salman et al., 2023) and a second 3-way exchange in India in December 2022, which was performed after our world-first 4-way exchange (Soin et al., 2023). Including the two 2-way exchanges performed before the computerized system was deployed in June 2022, the Malatya team had conducted a total of 15 liver exchange transplants by August 2023.
While the contribution of liver exchange to the number of LDLTs performed by the Malatya team was greater than that of any other liver exchange system worldwide up to that time, it still fell short of our expectations. For example, the 10 liver exchange transplants performed by the Malatya team in 2022 represented 4.3% of all LDLTs they carried out that year, whereas we had been hoping for at least 20%. Several factors account for this relatively slow start.
@@
Yılmaz must be one of the most perceptive and hardest-working individuals I have ever known. Despite being devastated by the situation, he navigated through this challenging period by collaborating with us to document our experience with the BBS–LPE System for scholarly publication, with a specific focus on our groundbreaking 4-way exchange.
@@
Fortunately, the editorial process in medicine moves much faster than in economics. The timely online publication of Yilmaz et al. (2023) in July 2023 not only provided an opportunity to introduce the BBS–LPE System to the general Turkish public but also reversed the unfortunate dynamics that had gripped the entire Malatya team.
12.6.4 From Pilot to Global Leadership: First Multi-Way Exchanges and Unprecedented Volume
Following the publication of Yilmaz et al. (2023), İnönü University leadership hosted a media event at the end of July 2023 to highlight the achievements of the BBS–LPE System. Participants from both the 4-way and 3-way exchanges attended, sharing their experiences and praise with the wider public. Emerging from a region still recovering from a once-in-a-century catastrophe, this success quickly captured national attention. Within weeks, patient participation surged, marking the conclusion of the BBS–LPE System’s pilot phase and the beginning of a period of rapid expansion.
@@ 12.6.4 From Pilot to Global Leadership: First Multi-Way Exchanges and Unprecedented Volume
Over the next 27 months, the system facilitated 315 additional liver exchange transplants—averaging more than 11 per month—bringing the cumulative total to 330 through October 2025. This rate exceeded that of the most active liver-exchange programs worldwide by an order of magnitude. Before the BBS–LPE System’s introduction, only one center—Medanta in India—had performed as many as ten liver-exchange transplants in any calendar year.
Despite the devastating earthquake that disrupted life across the region for months, the BBS–LPE System enabled 64 liver-exchange transplants in 2023, accounting for 27.7% of the 231 LDLTs performed by the Malatya team that year. Building on this momentum, it facilitated 122 transplants in 2024—an unprecedented 45.2% of the 270 LDLTs performed at the Institute. Further scaling up the program in 2025, the team had already performed 132 transplants by the end of October, with numbers continuing to rise. When its earlier exchanges are included (two transplants in 2021 and ten in 2022), the team’s cumulative total reached 330 liver-exchange transplants through October 2025 (see Figure 33).
Figure 33: BBS–LPE system: decomposition of liver exchanges by patient compatibility with their donors, as of October 2025. Acronyms: LPE: liver paired exchange; GRWR: graft-to-recipient weight ratio; ABO inc: ABO incompatible.
For perspective, fewer than 250 liver exchange transplants were reported worldwide between 2002 and 2022, before the launch of the BBS–LPE System. Ushering in a series of unprecedented milestones, in just three years, the Malatya team surpassed that two-decade global total, establishing its leadership in transplant volume and complexity:
@@
  • World’s First Multi-Way Exchanges: Building on its world-first 4-way liver exchange in July 2022, they performed the first 5-way, 6-way, and 7-way liver exchanges in October 2023, January 2024, and July 2024, respectively (Figure 34).
@@
  • Record-Breaking Cumulative Volume: By March 2024, they became the first to surpass 100 cumulative liver exchange transplants, in January 2025 the first to surpass 200, and in September 2025 the first to surpass 300.
@@
  • Unmatched Annual Performance: In October 2024, they became the first to exceed 100 liver exchange transplants in a single calendar year—a milestone repeated in September 2025.
Figure 34: Illustration of the world’s first 7-way liver exchange, performed by the BBS–LPE System at the Liver Transplantation Institute, İnönü University, Malatya, Turkey, on July 2, 2024. The exchange involved two ABO-incompatible pairs, two adult pairs with undersized grafts (GRWR << 0.8%), one pediatric pair with an oversized graft (GRWR >> 4.0%), and two compatible pairs who received more favorable ABO-identical grafts through exchange. Abbreviations: RL: Right lobe, ABO-id: ABO identical, ABO-c: ABO compatible but not identical, GRWR: Graft-to-recipient weight ratio.
As of October 2025, the BBS–LPE System accounts for both 7-way exchanges ever performed worldwide, all seven 6-way exchanges, all five 5-way exchanges, all thirteen 4-way exchanges, and twenty-nine 3-way exchanges—representing more than 80% of all 3-way liver exchange transplants ever reported globally (see Figure 35).
Figure 35: BBS–LPE system: decomposition of liver exchanges by the number of patient–donor pairs per exchange, as of October 2025. For each $k$-way exchange, bar width is proportional to the number of patients in each exchange, and bar area is proportional to the total number of patients who received transplants via $k$-way exchanges. Acronym: LPE: Liver paired exchange.
@@ 12.6.5 Comparison with Other Prominent Liver Exchange Programs
Prior to the introduction of the BBS–LPE System, India was the most active country globally for liver exchange. According to Kute et al. (2025), 259 liver exchange transplants were performed in the country between January 2007 and March 2025 across nine centers. This corresponds to 1.8% of the 14,341 LDLTs conducted in India during this period.
Almost half of these transplants were performed at Delhi-based Medanta—the global leader in liver exchange prior to BBS–LPE—which carried out 123 liver exchange transplants, including three 3-way and fifty-seven 2-way exchanges. These accounted for 2.6% of the 4,799 liver transplants performed at the center as of March 2025 (Soin et al., 2023; Kute et al., 2025). Another Delhi-based center, Max Saket Hospital, performed 76 liver exchange transplants in thirty-eight 2-way exchanges as of March 2025, constituting 1.6% of the 4,773 total liver transplants at that center (Agrawal et al., 2022; Kute et al., 2025).153153153Although Kute et al. (2025) does not disclose center names, Soin et al. (2023) for Medanta and Agrawal et al. (2022) for Max Saket Hospital—further verified through personal communication with Dr. Dhiraj Agarwal regarding the latter—confirm that these two centers are the highest- and second-highest-volume centers in India for kidney exchange.
Beyond India, another historically important program is at ASAN Medical Center in South Korea, a pioneer in kidney exchange. Liver exchange activity at the center and across the country has slowed in recent years with the expansion of ABO-incompatible liver transplantation. Although the ASAN team performed the world’s first liver exchange in 2002, only 26 liver exchange transplants were conducted between January 2003 and December 2011, corresponding to 1.2% of the 2,182 LDLTs performed at the center during that period (Jung et al., 2014). Nationwide, just 52 liver exchange transplants were reported in South Korea from 2002 to 2018, representing only 0.4% of the country’s LDLTs (Kim, 2022).
@@
Notably, all liver exchanges at Max Saket Hospital, and the vast majority at ASAN Medical Center, were limited to 2-way exchanges between blood-type incompatible pairs: one pair with a blood-type A patient and B donor, and another with a blood-type B patient and A donor. This restriction is likely the most significant factor explaining the relatively small number of exchanges at these centers.
@@
By contrast, in the U.S., liver exchange programs remain in their early stages. As of October 2025, a total of 196 liver exchange transplants had been reported nationwide, with a peak of 49 in 2023 (OPTN, 2025). Most activity has been concentrated at two centers. The program at the University of Texas Health San Antonio reported 23 liver exchange transplants from 2019 through October 2023, conducted through four 2-way exchanges, six 2-way non-directed donor chains, and one 3-way non-directed donor chain (Klair et al., 2024). Similarly, the program at the University of Pittsburgh Medical Center reported 20 liver exchange transplants between January 2019 and July 2021, carried out through three 2-way exchanges and seven 2-way non-directed donor chains (Gunabushanam et al., 2022).
In Turkey, liver exchange predates our establishment of the BBS–LPE system at İnöü University’s Liver Transplantation Institute. Dinckan et al. (2025) reports that 18 liver exchange transplants were performed in nine 2-way exchanges at İstinye University Hospital Multiorgan Transplant Center between January 2015 and December 2022. This represents one of the earliest large-scale liver exchange initiatives outside India and South Korea. Table 7 summarizes the major liver exchange programs worldwide that have reported 10 or more cumulative transplants.
Table 7: Major liver exchange programs with 10 or more cumulative transplants. For each country, the first row reports the national total, followed by rows for individual programs.
Country Program / Operator Start Documented volume
Turkey Multiple single-center programs 2015 $\geq$348   (by 10/2025)
İnönü University, Liver Transplantation Institute 2022 330   (by 10/2025)
İstinye University Hospital 2015 18   (01/2015–12/2022)
India Multiple single-center programs 2007 $\geq$259   (by 03/2025)
Medanta 2009 129   (by 03/2025)
Max Saket Hospital 2012 78   (2012–03/2025)
United States Multiple single-center programs 2014 196   (by 10/2025)
University of Texas Health San Antonio 2019 23   (01/2019-10/2023)
University of Pittsburgh Medical Center 2019 20   (01/2019–07/2021)
South Korea Multiple single-center programs 2002 $\geq$52   (2002–2018)
ASAN Medical Center 2003 26   (01/2003–12/2011)